OpenStudy (anonymous):

In a system of Cartesian coordinates are given points A = (0, 0) and P = (3, h). What's the alternative which its expression represents the distance from point P to point A as a function of h? a) d = 9 + h b) d= 9 + h² c) d = 9 + 6h + h² d) d = h + 3 Heeelp!

OpenStudy (bahrom7893):

d^2 = sqrt( (0-3)^2 + (0-h)^2 )

OpenStudy (bahrom7893):

I meant d = sqrt of all that

OpenStudy (anonymous):

umm d =sqrt(9+h^2)

OpenStudy (bahrom7893):

d = sqrt(9 + h^2)

OpenStudy (anonymous):

your options are wrong

OpenStudy (anonymous):

weird... I only have these options :/ why sqrt guys?

OpenStudy (anonymous):

apply pythagorus theorem, to undestand the formula correctly

OpenStudy (bahrom7893):